OpenStudy (anonymous):

The numbers on two consequtively numbered gym lockershave the sum of 127. What are the locker numbers?

OpenStudy (nikato):

let one of the number be x. the next consecutive number would be x+1 and since you know that the sum of these two consecutive numbers add up to 124, you can set up this equation x+x+1=127 Does this make sense so far?

OpenStudy (anonymous):

yes it does

OpenStudy (nikato):

so can you solve fro x on your own now?

OpenStudy (anonymous):

well do i have to find out two numbers that add up to 127 right?

OpenStudy (nikato):

correct. and you will be able to find these two numbers after you solve for x from that equation

OpenStudy (anonymous):

how would i solve for x if their is two x's

OpenStudy (nikato):

combine like terms remember x is the same as 1x so what's 1x+1x?

OpenStudy (anonymous):

2x

OpenStudy (nikato):

yes.so x+x+1=127 combining like terms give you 2x+1=127 what would be your next step to isolate x on the left side?

OpenStudy (anonymous):

i got 63

OpenStudy (anonymous):

it was 63,64 thanks for your help :)

OpenStudy (nikato):

Yes. That's correct. And you're welcome!